# Never too Cocky to Cooperate: An FIM and RL-based USV-AUV Collaborative System for Underwater Tasks in Extreme Sea Conditions

### v1

@@ III-B FIM-based USV Path Planning for AUV Positioning
13:           Update critics using $y\!=r+\gamma\min_{i=1,2}Q_{\theta_{i}^{\prime}}(s^{\prime},\tilde{a})$
@@
For the convenience of research without loss of generality, this work only considers the influence of different angles between AUVs of underwater targets on positioning accuracy. When the distance between the AUVs and the USV is equal and the AUVs are at the same depth, only the angle of the AUVs is considered, and the influence of AUV formation on positioning accuracy is only related to ${\rm\!sin}^{2}\alpha_{ij}$:
where ${\rm sin}\gamma_{0}\!\!=\!\!\frac{z_{k}}{S_{0}}$, while $\chi\!\!=\!\!\sum_{1\leq i\textless j\leq m}^{m}{\rm\!sin}^{2}\alpha_{ij}$, with $\alpha_{ij}\!\!=\!\!\varphi_{i}\!-\!\varphi_{j}$ representing the angle between the projections of two AUVs and USV.
Through calculating the waypoints using Eq. (8) with the maximum $\rm det\left(\mathbf{{\textbf{{J}}}}_{m}\right)$, we can implement USV path planning to minimize AUV positioning errors through above process.
III-C RL-Driven Multi-AUV Decision-Making
Our approach leverages the Twin Delayed Deep Deterministic Policy Gradient (TD3) algorithm, which extends the Deep Deterministic Policy Gradient (DDPG) algorithm with three key improvements: (1) clipped double Q-learning to mitigate overestimation bias, (2) delayed policy updates to stabilize training, and (3) target policy smoothing to prevent overfitting to narrow peaks in the value function [36]. These characteristics make TD3 particularly suitable for underwater vehicle control, where precise continuous actions are required and the dynamics are highly coupled across degrees of freedom.
$<\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{P}(\cdot|s,a),\gamma>,$ (9)
where continuous state space $\mathcal{S}$, continuous action space $\mathcal{A}$, reward function $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$. $\mathcal{P}(\cdot|s,a)\sim(0,1)$ is the state transition probability. $\gamma\in[0,1]$ represents the discount factor.
Given state space $\mathcal{S}$ and action space $\mathcal{A}$, our objective in the underwater tasks is to obtain the expert policy through maximizing discounted cumulative reward:
where $\pi:\mathcal{S}\to\mathcal{A}$ is a deterministic policy. The agent’s objective is to find the optimal policy $\pi^{*}$ that maximizes the expected return: where $\gamma\in[0,1)$ is the discount factor. This infinite-horizon formulation requires careful treatment of the value function approximation.
The target noise $\epsilon$ serves as a regularizer, preventing the policy from exploiting overestimated Q-values. The clipped Gaussian noise ensures the perturbation stays within reasonable bounds:
@@ III-C RL-Driven Multi-AUV Decision-Making
The target value computation uses the minimum of these two estimates: Maintain two Q-networks $Q_{\theta_{1}}$ and $Q_{\theta_{2}}$ with target networks:
@@
Thus, we can obtain Critic’s loss function is based on the TD error of the Bellman equation:
@@
The policy network updates less frequently than the Q-functions (typically every $d=2$ steps). This delay allows the value estimates to stabilize before the policy is adjusted, reducing error propagation. According to the deterministic policy gradient theorem
@@
we can get the policy gradient, which is computed as:
@@
Note that only $Q_{\theta_{1}}$ is used for the policy gradient to prevent the ”chasing” behavior that could occur if both Q-networks were used simultaneously.
In practice, we can estimate the policy gradient using Monte Carlo method:
@@
Finally, we update the policy network, and update the target networks using soft update strategy. The overall RL training using TD3 algorithm has been shown in pseudo-code Algorithm 1.
(a) Sea conditions at 25s
(b) Sea conditions at 50s
(c) Sea conditions at 75s
@@ IV Experiments and Analysis
TABLE I: Parameters of the Environment and Algorithm
@@
Hidden layer size $128$
@@
Target policy noise $\mathcal{N}(0,0.1)$, clipped to ±$\pm 1.0$
@@ IV-A Task Description and Experimental Settings
Given the limited availability of open-source underwater tasks, we selected the underwater data collection task as a benchmark scenario to validate our system. In this application scenario, multiple AUVs and a USV work collaboratively to collect data from sensor nodes (SNs) deployed in the Internet of Underwater Things (IoUT) environments, while simultaneously optimizing multiple objectives including seviced SN number (SSN) and sum data rate (SDR) maximization, collision avoidance, and energy consumption (EC) minimization. The specific parameter configurations employed in our study are comprehensively listed in Table 1. For further technical specifications and detailed task parameters, readers may consult our prior research work [5].
@@
The simulation experiments are conducted on a personal computer equipped with an AMD Ryzen 9 5950X processor and an NVIDIA GeForce RTX 4060 GPU, operating within a Python 3.8 virtual environment. Due to the adoption of an efficient computational approach for tidewave simulation, our code required approximately 3 hours to complete 600 episodes under the aforementioned configuration, demonstrating low computational resource demands. For other parameters, please refer to Table 1 for summary.
@@ IV-B Realistic Extreme Sea Condition Simulation
The USV-AUV collaborative system’s dual-domain operation (surface and subsurface) makes it particularly vulnerable to marine disturbances. To address this, we employ two-dimensional tidal wave equations and ocean turbulence models to realistically simulate extreme sea conditions, accurately replicating disturbances that impact AUV positioning and performance for rigorous system validation. The simulation effects of sea conditions at three different time points (25s, 50s, 75s) have been shown in Fig. 3.
TABLE II: Comparison of Different RL Algorithms in Ideal and Extreme Sea Conditions
SDR EC ARPS SSN
TD3 (ISC) 11.20±1.93 138.19±3.80 -60.69±7.81 47.0±5.52
TD3 (ESC) 10.78±1.70 147.81±5.09 -62.95±8.60 45.4±6.81
DDPG (ISC) 6.63±0.97 73.72±1.11 -80.70±6.94 30.8±3.97
DDPG (ESC) 6.35±1.57 77.23±0.89 -82.28±7.41 29.6±4.91
SAC (ISC) 5.97±2.66 111.57±13.23 -87.86±13.10 27.2±7.35
SAC (ESC) 5.58±2.35 117.76±17.30 -89.58±15.22 25.0±6.62
Baseline [5] 5.93±3.51 112.30±10.56 -87.95±12.08 25.8±7.94
If we denote the wave velocity as $\bm{V}\!_{w}\!=\![u,v]$, and the gravitational acceleration as $g$, the water level $\eta$ at coordinate point $(x^{{}^{\prime}},y^{{}^{\prime}}\!)$ can be calculated using Finite Difference Method (FDM). The equations’ forward difference scheme using FDM can be listed as follows [37]:
@@
(a) Average reward per timestep
(b) Sum data rate
(c) Energy consumption
(d) Serviced Sensor Nodes
@@ IV-C State, Action Space and Reward Function Design
where $\{{\Delta\mathbf{p}_{j}^{k}\}_{j\neq k}}=\frac{\mathbf{p}_{t}^{j}-\mathbf{p}_{t}^{k}}{\|\mathbf{b}\|}$ denotes relative positions of other AUVs, with $\mathbf{p}_{t}^{k}\in\mathbb{R}^{2}$ denoting the 2D position of the $k$-th AUV, and $\mathbf{b}=[X_{\text{max}},Y_{\text{max}}]^{\top}$ defining the environment boundaries. Besides, $\Delta\mathbf{p}_{\text{target}}^{k}=\frac{\mathbf{p}_{\text{target}}^{k}-\mathbf{p}_{t}^{k}}{\|\mathbf{b}\|}$ denotes the relative position to target SN, with $\mathbf{p}_{\text{target}}^{k}$ representing the position of the target SN assigned to the $k$-th AUV. Normalized self-position $\tilde{\mathbf{p}}_{t}^{k}=\frac{\mathbf{p}_{t}^{k}}{\|\mathbf{b}\|}$, while $\rho_{\text{overflow}}=\frac{N_{\text{DO}}}{N_{\text{POI}}}$ is the data overflow ratio. Additionally, $\mathbb{I}_{\text{border}}^{k}\in\{0,1\}$ represents the border violation flag.
The full action vector of action $a_{t}^{k}\in\mathcal{A}$ for the $k$-th AUV is a 2D continuous vector:
@@
where $v_{\text{norm}}$ denotes expected normalized speed, which is linearly mapped to the physical range $[V_{\text{min}},V_{\text{max}}]$: $v_{\text{norm}}=\frac{2(v_{t}^{k}-V_{\text{min}})}{V_{\text{max}}-V_{\text{min}}}-1$. $\theta_{\text{norm}}$ stands for expected normalized direction, and is scaled to $[-\pi,\pi]$: $\theta_{\text{norm}}=\frac{\theta_{t}^{k}}{\pi}$.
Since reward is significant for RL training, we employ a shaped reward function that balances multiple objectives:
@@
where $\bm{R}=\{R_{1},R_{2},\ldots,R_{p}\}$ represents distinct objectives. $\bm{\lambda}=\{\lambda_{1},\lambda_{2},\ldots,\lambda_{p}\}$ control the relative weighting of these objectives. Substituting into underwater data collection task, our reward function is designed as follows:
$\begin{split}\mathcal{R}=-0.6d_{k}^{\text{target}}&-0.1\mathbb{I}_{\text{border}}-0.05N_{\text{DO}}-\sum_{j\neq k}6(12-d_{kj})\\ &+12\mathbb{I}_{\text{TL}}-0.085E_{k},\end{split}$ (24)
specifically, the terms $d_{k}^{\text{target}}$ denotes AUV-SN distance, $\mathbb{I}_{\text{border}}$ represents border violation indicator, while $N_{\text{DO}}$ stands for data overflow count. Besides, $d_{kj}$ is inter-AUV distance, and $\mathbb{I}_{\text{TL}}$ denotes transmission success flag, while $E_{k}$ represents energy consumption.
@@ IV-D Experiment Results and Analysis
To comprehensively assess the feasibility and effectiveness of our proposed system, we conducted a series of systematic experiments involving both training and evaluation phases. Initially, we implemented three state-of-the-art RL algorithms, TD3, DDPG and Soft Actor-Critic (SAC), to train two AUVs that were positioned by a USV. The training process was carried out under two distinct marine environments: ideal sea conditions and extreme sea conditions. As demonstrated in Fig. 4, the learning curves exhibit stable convergence and eventually reach performance levels comparable to expert policies, clearly indicating that the AUVs have successfully learned optimal navigation and data collection strategies through the RL training process. While we can also find that TD3 outperforms other algorithms and baseline in SDR and ARPT, although energy consumption is higher than others.
@@
In addition, to further validate the system’s robustness, we conducted extensive environmental generalization tests comparing system performance between ideal and extreme sea conditions. The experimental results, summarized in Table 2 (where ISC and ESC represent Ideal Sea Conditions and Extreme Sea Conditions, while ARPS and SSN denote Average Reward Per Step and Serviced Sensor Nodes, respectively), reveal that despite the challenging environmental factors present in extreme conditions, the system maintains comparable performance metrics to those achieved in ideal conditions. This remarkable consistency in performance across different environmental scenarios strongly demonstrates the system’s exceptional robustness and adaptability when operating in extreme sea conditions.
@@
Figure 6: Negative determinant value evolution of FIM in an operational episode using different number of AUVs. (a) 2 AUVs. (b) 3 AUVs. (c) 4 AUVs.
@@
Figure 7: The illustrations of the USV-AUV coordinate relationships in the collaborative system at some point. The optimal USV positions were determined through real-time optimization of FIM determinant value to minimize AUV positioning errors. (a) 2 AUVs: The optimal USV position is at coordinates [61.3, 26.5]. (b) 3 AUVs: The optimal USV position is at coordinates [45.2, 31.1]. (c) 4 AUVs: The optimal USV position is at coordinates [39.8, 39.2].
To further validate the scalability of the proposed USV-AUV collaboration framework, we conducted a new set of experiments using 1 AUV, 2 AUVs, 3 AUVs and 4 AUVs for underwater data collection tasks, evaluating their performance in extreme sea conditions, respectively. The results were plotted in the form of a histogram in Fig. 5. From the histogram we can obtain that as the number of AUVs increases, although ARPT shows a downward trend, the system’s SDR improves significantly. Specifically, when the number of AUVs increases from 1 to 4, the SDR and SSN increases by approximately 14.7 and 62, respectively. This indicates that multi-AUV collaboration can effectively distribute task loads and significantly enhance overall efficiency through parallel data collection. Although the increase in the number of AUVs also leads to higher EC, the growth remains within a reasonable range. For example, the EC for 4 AUVs is only about 1% higher than that for 3 AUVs and even 10% lower for a single AUV, which is far below the proportional improvement in SDR. This demonstrates the efficiency of the proposed system in optimizing resource allocation and path planning, ensuring performance gains while keeping additional energy consumption at an acceptable level.
@@
Figure 9: Positioning error of the AUV with USV fixed at (0,0) and (100, 100), and path planning using FIM optimization, respectively.
@@
To further evaluate the robustness and scalability of the proposed USV-AUV collaborative system, we conducted additional experiments employing 2, 3, and 4 AUVs for underwater data collection tasks in extreme sea conditions. A key focus of these experiments was to minimize the negative determinant of the FIM in real-time during each operational episode, thereby optimizing USV path planning for enhanced AUV positioning accuracy. The results demonstrate that despite significant environmental disturbances (including ocean turbulence and strong waves), the system consistently maintained the FIM determinant’s negative value within a minimized range throughout each operational episode. As shown in Fig. 6, this achievement was observed across all tested configurations (2, 3, and 4 AUVs), with the minimization performance showing particular robustness in the extreme sea conditions. Furthermore, through visualization in Fig. 7, it was clearly observed that the USV’s path planning points consistently fell in the region with the lowest AUV positioning error. The optimal USV positions in the USV-AUV collaborative system were determined as follows:
  • 2 AUVs configuration: Optimal USV position at [61.3, 26.5] for AUVs located at [25, 93] and [95, 40].
  • 3 AUVs configuration: Optimal position at [45.2, 31.1] for AUVs at [20, 90], [95, 45], and [30, 14].
  • 4 AUVs configuration: Optimal position at [39.8, 39.2] for AUVs at [36, 93], [55, 44], [15, 19], and [72, 26].

### v2

@@ III-C RL-Driven Multi-AUV Decision-Making
[…]
@@
The target noise $\epsilon$ serves as a regularizer, preventing the policy from exploiting overestimated Q-values. The clipped Gaussian noise ensures that the perturbation stays within reasonable bounds:
@@
The target value computation uses the minimum of these two estimates. Specifically, we maintain two Q-networks $Q_{\theta_{1}}$ and $Q_{\theta_{2}}$ with target networks:
@@
Thus, we can obtain that the Critic’s loss function is based on the TD error of the Bellman equation:
@@
The policy network updates less frequently than the Q-functions (typically every $d=2$ steps). This delay allows the value estimates to stabilize before the policy is adjusted, thereby reducing error propagation. According to the deterministic policy gradient theorem
@@
we can obtain the policy gradient, which is computed as:
@@
Note that only $Q_{\theta_{1}}$ is used for the policy gradient to prevent the ”chasing” behavior that could occur if both Q-networks are used simultaneously.
In practice, we can estimate the policy gradient using the Monte Carlo method:
@@
Finally, we update the policy network, and update the target networks using soft update strategy. The overall RL training using the TD3 algorithm has been shown in pseudo-code Algorithm 1.
@@ IV Experiments and Analysis
TABLE II: Parameters of the Environment and Algorithm
@@
Hidden layer size 128128
@@
Target policy noise $\mathcal{N}(0,0.1)$, clipped to $\pm 1.0$
@@ IV-A Task Description and Experimental Settings
Given the limited availability of open-source underwater tasks, we selected the underwater data collection task as a benchmark scenario to validate our system. In this application scenario, multiple AUVs and a USV work collaboratively to collect data from sensor nodes (SNs) deployed in the Internet of Underwater Things (IoUT) environments while simultaneously optimizing multiple objectives, including the serviced SN number (SSN) and sum data rate (SDR) maximization, collision avoidance, and energy consumption (EC) minimization. The specific parameter configurations employed in our study are comprehensively listed in Table I. For further technical specifications and detailed task parameters, readers may consult our prior research work [5].
@@
The simulation experiments are conducted on a personal computer equipped with an AMD Ryzen 9 5950X processor and an NVIDIA GeForce RTX 4060 GPU, operating within a Python 3.8 virtual environment. Due to the adoption of an efficient computational approach for tidal wave simulation, our code required approximately 3 hours to complete 450 episodes under the aforementioned configuration, demonstrating low computational resource demands. For other parameters, please refer to Table II for a summary.
@@ IV-B Realistic Extreme Sea Condition Simulation
The USV-AUV collaborative system’s dual-domain operation (surface and subsurface) makes it particularly vulnerable to marine disturbances. To address this, we employ two-dimensional tidal wave equations and ocean turbulence models to realistically simulate extreme sea conditions, accurately replicating disturbances that impact AUV positioning and performance for rigorous system validation. The simulation effects of sea conditions at three different time points (25 s, 50 s, 75 s) are shown in Fig. 3.
If we denote the wave velocity as $\bm{V}\!_{w}\!=\![u,v]$, and the gravitational acceleration as $g$, the water level $\eta$ at coordinate point $(x^{{}^{\prime}},y^{{}^{\prime}}\!)$ can be calculated using the Finite Difference Method (FDM). The equations’ forward difference scheme using FDM can be listed as follows [36]:
@@ IV-C State, Action Space and Reward Function Design
where $\{{\Delta\mathbf{p}_{j}^{k}\}_{j\neq k}}=\frac{\mathbf{p}_{t}^{j}-\mathbf{p}_{t}^{k}}{\|\mathbf{b}\|}$ denotes the relative positions of other AUVs, with $\mathbf{p}_{t}^{k}\in\mathbb{R}^{2}$ denoting the 2D position of the $k$-th AUV, and $\mathbf{b}=[X_{\text{max}},Y_{\text{max}}]^{\top}$ defining the environmental boundaries. Besides, $\Delta\mathbf{p}_{\text{target}}^{k}=\frac{\mathbf{p}_{\text{target}}^{k}-\mathbf{p}_{t}^{k}}{\|\mathbf{b}\|}$ denotes the relative position to the target SN, with $\mathbf{p}_{\text{target}}^{k}$ representing the position of the target SN assigned to the $k$-th AUV. The normalized self-position $\tilde{\mathbf{p}}_{t}^{k}=\frac{\mathbf{p}_{t}^{k}}{\|\mathbf{b}\|}$ encodes spatial awareness within the bounded domain. Moreover, $\rho_{\text{overflow}}=\frac{N_{\text{DO}}}{N_{\text{POI}}}$ represents the data overflow ratio, a global congestion indicator reflecting the proportion of sensor nodes whose data buffers are full at time $t$. A higher $\rho_{\text{overflow}}$ implies severe network congestion and guides AUVs to prioritize heavily loaded nodes for data collection. In contrast, $\mathbb{I}_{\text{border}}^{k}\in\{0,1\}$ is the border violation flag, which takes the value 11 when the predicted next position $\tilde{\mathbf{p}}_{t+1}^{k}=\mathbf{p}_{t}^{k}+v_{t}^{k}\Delta t[\cos(\theta_{t}^{k}),\sin(\theta_{t}^{k})]^{\top}$ lies outside the operational boundary $\mathcal{Q}=[0,X_{\max}]\times[0,Y_{\max}]$. Including both quantities in the state vector allows the agent to simultaneously perceive global network congestion and local safety constraints, which are essential for maintaining balanced and efficient underwater operations.
The full action vector $a_{t}^{k}\in\mathcal{A}$ for the $k$-th AUV is a 2D continuous vector:
@@
where $v_{\text{norm}}$ denotes the expected normalized speed, which is linearly mapped to the physical range $V_{\text{min}},V_{\text{max}}]$: $v_{\text{norm}}=\frac{2(v_{t}^{k}-V_{\text{min}})}{V_{\text{max}}-V_{\text{min}}}-1$. $\theta_{\text{norm}}$ stands for the expected normalized direction and is scaled to $\pi,\pi]$: $\theta_{\text{norm}}=\frac{\theta_{t}^{k}}{\pi}$.
(a) Sea conditions at 25s
(b) Sea conditions at 50s
(c) Sea conditions at 75s
TABLE III: Comparison of Different RL Algorithms in Ideal and Extreme Sea Conditions
SDR (Mbps) EC (W) ARPT SSN Trajectory Lengths (m)
TD3 (ISC) 11.20±1.93 138.19±3.80 -60.69±7.81 47.0±5.52 (1354±75,1340±45)
TD3 (ESC) 10.78±1.70 147.81±5.09 -62.95±8.60 45.4±6.81 (1389±38,1371±34)
DDPG (ISC) 6.63±0.97 73.72±1.11 -80.70±6.94 30.8±3.97 (1138±41,1089±63)
DDPG (ESC) 6.35±1.57 77.23±0.89 -82.28±7.41 29.6±4.91 (1108±34,1095±42)
SAC (ISC) 5.97±2.66 111.57±13.23 -87.86±13.10 27.2±7.35 (1156±89,1171±78)
SAC (ESC) 5.58±2.35 117.76±17.30 -89.58±15.22 25.0±6.62 (1132±94,1109±62)
Baseline [5] 5.93±3.51 112.30±10.56 -87.95±12.08 25.8±7.94 (1170±71,1196±77)
Since the reward plays a critical role in RL training, we design a shaped reward function that balances multiple objectives:
@@
where $\bm{R}=\{R_{1},R_{2},\ldots,R_{p}\}$ represents distinct objectives, and $\bm{\lambda}=\{\lambda_{1},\lambda_{2},\ldots,\lambda_{p}\}$ controls the relative weighting of these objectives.
Substituting into the underwater data collection task, our reward function is designed as follows:
$\begin{split}\mathcal{R}=\underbrace{-0.6d_{k}^{\text{target}}-0.05N_{\text{DO}}+12\mathbb{I}_{\text{TL}}}_{\text{Approaching target SNs}}-\underbrace{0.085E_{k}}_{\text{Energy consumption}}\\ \underbrace{-\sum_{j\neq k}6(12-d_{kj})-0.1\mathbb{I}_{\text{border}}}_{\text{Safety requirements}},\end{split}$ (24)
where $d_{k}^{\text{target}}$ denotes the AUV–SN distance, $\mathbb{I}_{\text{border}}$ and $N_{\text{DO}}$ correspond to the border violation and network congestion penalties, $d_{kj}$ is the inter-AUV distance, $\mathbb{I}_{\text{TL}}$ represents the transmission success flag, and $E_{k}$ indicates energy consumption. Accordingly, $N_{\text{DO}}$ and $\mathbb{I}_{\text{border}}^{k}$ jointly influence the RL mechanism: the former informs the policy regarding the overall data backlog, while the latter constrains unsafe maneuvers near domain boundaries. Together, they allow the policy to learn a cooperative behavior that enhances both communication throughput and navigation safety.
(a) Sum data rate
(b) Energy consumption
(c) Serviced sensor nodes
Figure 5: Performance comparison under different weighting ratios between energy- and SN-related reward terms. The results illustrate the trade-off between energy efficiency and data collection capability when scaling the relative weights of these two terms. (a) Sum data rate. (b) Energy consumption. (c) Serviced sensor nodes.
(a) Safety violation times
(b) Sum data rate
(c) Energy consumption.
(d) Serviced Sensor Nodes.
Figure 6: Performance trends under different safety-related reward coefficients. The plots show the sensitivity of key performance metrics to the scaling of the safety weight. (a) Safety violation times. (b) Sum data rate. (c) Energy consumption. (d) Serviced Sensor Nodes.
@@ IV-D Experiment Results and Analysis
To comprehensively evaluate the feasibility and effectiveness of the proposed system, we conducted a series of systematic experiments involving both training and evaluation phases, where three state-of-the-art RL algorithms: TD3, DDPG, and Soft Actor-Critic (SAC) were implemented to train two AUVs coordinated by a USV under both ideal and extreme sea conditions. As shown in Fig. 4, the RL training process exhibits stable convergence across random seeds, and the system ultimately reaches an expert-level policy, indicating consistent and robust learning behavior. Here, the convergence is declared only when the 25-episode moving averages of these metrics exhibit slope magnitudes below 0.2 for at least 50 consecutive episodes and no further improvement is observed in validation metrics, ensuring that the selected checkpoint corresponds to a fully converged rather than prematurely stopped policy. Therefore, the expert policy here refers to the fully converged model that achieves the highest validation performance across these metrics. Quantitatively, the Sum Data Rate (SDR) of TD3 increases rapidly and stabilizes around 11 Mbps, approximately 70% higher than that of DDPG and SAC ($\approx 6$ Mbps) under ideal conditions. In contrast, although TD3 incurs a slightly higher peak Energy Consumption (EC), which is roughly 20% higher than SAC, it maintains the highest Average Reward Per Timestep (ARPT) ($\approx-60$), showing a 25–30% improvement over both DDPG and SAC in long-term task performance. Taken together, these quantitative comparisons confirm that TD3 not only converges stably to a rigorously defined expert policy but also achieves the best trade-off between data collection efficiency and energy cost, significantly outperforming DDPG and SAC across key evaluation metrics.
@@
In addition, to further assess the robustness of the proposed system, we conducted extensive environmental generalization experiments comparing performance under both Ideal Sea Conditions (ISC) and Extreme Sea Conditions (ESC). The results, summarized in Table III (where SSN denotes Serviced Sensor Nodes) show that, despite the challenging environmental factors in ESC, the system maintains performance levels comparable to those achieved under ISC. This consistent performance across different environmental scenarios highlights the system’s strong robustness and adaptability in extreme sea conditions.
Besides, we further introduced the “total distance traveled per AUV” (denoted as trajectory lengths in the tables) as an additional performance metric to complement these evaluations and provide a more comprehensive understanding of cooperative behavior and task assignment. As shown in Table III, the trajectory lengths of individual AUVs under both ISC and ESC remain highly consistent, for example, (1354 ± 75 m, 1340 ± 45 m) for TD3 (ISC) and (1389 ± 38 m, 1371 ± 34 m) for TD3 (ESC), with a variance below 2%. This quantitative stability demonstrates that the proposed collaborative system effectively balances workload distribution among AUVs, preventing any single vehicle from being overburdened. Moreover, compared to baseline methods, the utilized approach (TD3) achieves an average increase of about 80% in served SNs with only approximately an 20% increase in trajectory lengths, confirming the system’s superiority in path-planning efficiency and energy conservation.
(a) Average reward per timestep
(b) Sum data rate
(c) Energy consumption
(d) Serviced Sensor Nodes
TABLE IV: Comprehensive Performance Comparison under Different Numbers of AUVs
SDR (Mbps) EC (W) ARPT SSN Trajectory Lengths (m)
$N=1$ 5.94±1.03 111.47±5.79 -23.87±3.27 26.8±1.04 (1309±91)
$N=2$ 10.78±1.70 147.81±5.09 -62.95±8.60 45.4±6.81 (1389±38,1371±34)
$N=3$ 16.39±2.90 139.72±4.18 -78.26±7.84 70.3±12.45 (1418±108,1452±72,1380±71)
$N=4$ 21.02±2.94 124.83±6.63 -92.21±7.51 88.7±13.74 (1445±125,1407±159,1512±153,1420±86)
Building upon the above experimental evaluations, we further delve into the design of the reward function and the determination of a robust set of reward weights. As shown in Eq. 24, the reward integrates interpretable terms for target-tracking ($d_{k}^{\text{target}}$), safety ($\mathbb{I}_{\text{border}}$ and $d_{kj}$), data overflow and transmission success ($N_{\text{DO}}$ and $\mathbb{I}_{\text{TL}}$), and energy consumption ($E_{k}$). Correspondingly, we analyzed how the relative weighting of these factors affects system behavior. Since safety is fundamental for multi-AUV coordination, safety-related terms were prioritized, which effectively reduced boundary violations and collision risks to negligible levels. After ensuring stable safe behaviors, we examined the trade-off between data-collection efficiency (affected by $d_{k}^{\text{target}}$, $N_{\text{DO}}$, $\mathbb{I}_{\text{TL}}$) and energy usage ($E_{k}$), revealing a Pareto relationship between operational gain and energy expenditure.
@@
Figure 8: Negative determinant value evolution of FIM in an operational episode using different numbers of AUVs. (a) 2 AUVs. (b) 3 AUVs. (c) 4 AUVs.
@@
Figure 9: The illustrations of the USV-AUV coordinate relationships in the collaborative system at some point. The optimal USV positions were determined through real-time optimization of FIM determinant value to minimize AUV positioning errors. (a) 2 AUVs: The optimal USV position is at coordinates (28.67 m, 46.01 m). (b) 3 AUVs: The optimal USV position is at coordinates (39.50 m, 33.45 m). (c) 4 AUVs: The optimal USV position is at coordinates (43.77 m, 40.17 m).
@@
Figure 11: Positioning error of the AUVs with USV fixed at (0,0) and (100, 100), and path planning using FIM optimization, respectively.
TABLE V: Comparison of Major Performance Metrics under Different USV Path Planning Policies
SDR (Mbps) EC (W) ARPT SSN
TD3 (ESC) 10.78±1.70 147.81±5.09 -62.95±8.60 45.4±6.81
TD3-R/S (ESC) 9.28±2.74 154.14±9.40 -65.89±9.45 39.0±10.35
TD3-R/D (ESC) 10.47±2.23 150.37±9.51 -64.38±8.25 44.9±8.41
TABLE VI: Comparison of Major Performance Metrics under Varying Numbers of FIM Optimization Iterations
Iteration counts SDR (Mbps) EC (W) ARPT SSN
nit=6 8.67±2.29 156.31±9.23 -67.51±10.32 37.3±7.93
nit=12 10.08±1.66 155.45±6.25 -65.64±6.30 43.0±6.82
nit=18 10.40±1.88 149.92±9.11 -63.81±7.56 44.5±6.68
nit=24 10.82±2.10 152.17±7.56 -62.09±8.81 46.3±7.57
(a) Positioning error of AUVs under different USV path planning policies
(b) Positioning error of AUVs under different FIM optimization iterations
Figure 12: AUV positioning error under different USV path planning policies and different number of FIM optimization iterations. (a) Positioning error of AUVs under different USV path planning policies. (b) Positioning error of AUVs under different FIM optimization iterations.
We first investigate how the relative weighting between the other two reward components influences task performance. To this end, we scaled their weights by factors of $(0.25,0.5,1,2)$. Controlled experimental results in Fig. 5 showed that overly small energy penalties led to high-speed but energy-intensive trajectories, whereas excessive penalization limited mobility and reduced the SSNs; moderate weighting achieved an optimal balance across the SDR, EC, and SSN. Subsequently, we multiply the safety-related term by scaling factors ranging from 0.1 to 10 and analyze how safety performance (i.e., the total violation times of all AUVs) and other major metrics evolve. As depicted in Fig. 6, further tuning of the safety coefficient demonstrated that increasing its weight rapidly suppressed violations near a ratio of one, beyond which improvements saturated and minor performance degradation appeared. Based on these observations, we selected a safety weight near this knee point and balanced SN-tracking and energy weights around unity, yielding the final configuration that ensures stable convergence and strong generalization in both ideal and extreme sea conditions.
Based on the above results, we further evaluated the scalability and coordination efficiency of the proposed USV–AUV collaborative system by testing configurations with 1, 2, 3, and 4 AUVs under extreme sea conditions. As summarized in Table IV and illustrated in Fig. 7, the system exhibits clear and consistent performance trends as the team size increases. Although ARPT shows a gradual decline due to higher system complexity, both SDR and SSN improve markedly: SDR increases from 5.9 Mbps to 21.0 Mbps, and SSN increases from 26.8 to 88.7 as the number of AUVs grows from 1 to 4. This pattern arises from the saturation of discrete positive rewards and the accumulation of energy, waiting, and proximity penalties that scale with fleet size. Hence, the declining ARPT reflects denser penalty events rather than degraded performance. Mission-level metrics (SDR and SSN) confirm that overall efficiency and scalability remain robust, demonstrating that multi-AUV cooperation effectively distributes task loads and enhances overall operational efficiency through parallelized data collection. Meanwhile, the trajectory lengths remain highly consistent among all AUVs across different configurations; for example, they are (1389 ± 38 m, 1371 ± 34 m) for $N=2$ and (1445 ± 125 m, 1407 ± 159 m, 1512 ± 153 m, 1420 ± 86 m) for $N=4$, indicating that the cooperative RL framework achieves balanced workload allocation without overburdening any single vehicle. Although larger teams naturally entail higher energy consumption, the increase remains moderate; for instance, the EC for four AUVs is only about 1% higher than that for three AUVs and even 10% lower than that for a single AUV, confirming that the proposed framework can scale efficiently, optimize path planning and resource utilization, and maintain robust, energy-aware coordination even under extreme sea conditions.
@@
To further evaluate the robustness and scalability of the proposed USV-AUV collaborative system, we conducted additional experiments employing 2, 3, and 4 AUVs for underwater data collection tasks in extreme sea conditions. A key focus of these experiments was to minimize the negative determinant of the FIM in real-time during each operational episode, thereby optimizing USV path planning for enhanced AUV positioning accuracy. The results demonstrate that, despite significant environmental disturbances (including ocean turbulence and strong waves), the system consistently maintained the FIM determinant’s negative value within a minimized range throughout each operational episode. As shown in Fig. 8, this achievement was observed across all tested configurations (2, 3, and 4 AUVs), with the minimization performance demonstrating particular robustness in extreme sea conditions. Furthermore, through visualization in Fig. 9, it was clearly observed that the USV’s path planning points consistently fell in the region with the lowest AUV positioning error. The optimal USV positions in the USV-AUV collaborative system were determined as follows:
  • 2 AUVs configuration: Optimal USV position at (28.67 m, 46.01 m) for AUVs located at (25 m, 93 m) and (95 m, 40 m).
  • 3 AUVs configuration: Optimal USV position at (39.50 m, 33.45 m) for AUVs located at (20 m, 90 m), (95 m, 45 m), and (30 m, 14 m).
  • 4 AUVs configuration: Optimal USV position at (43.77 m, 40.17 m) for AUVs located at (36 m, 93 m), (65 m, 54 m), (15 m, 19 m), and (92 m, 26 m).